# GRID: Visual Layout Generation

### v1

@@ Abstract
Visual Layout Generation
@@ 3.2 Main Results
AnimateDiffv3 464.1 35.24 12 - - - 250.9 0.9229
VideoCrafter2 424.2 32.00 15 - - - - -
OpenSora1.2 472.0 39.07 12 1000.5 37.11 66 - -
CogVideo5b 301.1 36.27 48 740.1 34.82 132 122.5 0.9185
Ours(Stage1) 482.1 32.46 7.2 1003.2 32.48 24 115.5 0.9598
@@
As shown in Table 1, we evaluate our approach on both short (16-frame) and long (64-frame) generation tasks. For 16-frame generation, our full model pushes FVD to 418.9, improving the previous best score by 15%, while achieving a 37.34 IS score. More notably, we significantly reduce computational costs, achieving 6.7× faster inference (7.2s vs 48s) compared to CogVideo. This efficiency advantage becomes more pronounced in 64-frame generation, where our approach maintains 5.5× faster inference speed. Our staged training demonstrates clear benefits: Stage1 (trained with coarse annotations and $\mathcal{L}_{base}$ only) achieves a FVD of 482.1, which reduces to 438.3 with fine-grained annotations, and further reduces to 418.9 with $\mathcal{L}_{flow}$.
@@ A.2 Why is it Natural for GRID to Leverage Built-in Attention Mechanism
A.3 Computational Efficiency Analysis
Current state-of-the-art video generation models require extensive computational resources. CogVideo utilizes approximately 35M video clips with an additional 2B filtered images from LAION-5B and COYO-700M datasets. Open-Sora operates on more than 30M videos and demands multiple 80GB GPUs with sequence parallelism for inference.
In contrast, GRID achieves competitive performance with less than 35K training videos and requires only 160M-2560M additional parameters. This significant reduction in resource requirements stems from our grid-based architecture and targeted training strategy. Specifically, our architecture enables efficient parallel processing and parameter sharing, while our training approach focuses on essential video-specific components while leveraging pretrained image capabilities.
The efficiency improvements can be quantified through the following key aspects:
@@ A.3 Computational Efficiency Analysis
$\text{Data Reduction}\approx\frac{>35M}{<35K}=1000\times$ (11)
@@
These improvements enable high-quality video generation while maintaining accessibility for research environments with limited computational resources.
A.4 Inference Details
@@
A.5 Post-Processing Pipeline
@@ A.6 Potential Applications
A.6.1 Creative Multi-view Generation
@@
A.6.2 Flexible Frame Extension
@@
A.6.3 Future Extension to Video Understanding
@@
A.6.4 Maintained Image Generation Ability
@@ A.7 Limitations
A.8 Multyview Camera Parameters

### v3

# GridShow: Omni Visual Generation

@@ Abstract
Visual Grid Generation
@@ 3.2 Main Results
AnimateDiffv3 464.1 35.24 12 - - - 250.9 0.9229 419M
VideoCrafter2 424.2 32.00 15 - - - - - 919M
OpenSora1.2 472.0 39.07 12 1000.5 37.11 66 - - 1.5B
CogVideo5b 301.1 36.27 48 740.1 34.82 132 122.5 0.9185 5B
Ours(Stage1) 482.1 32.46 7.2 1003.2 32.48 24 115.5 0.9598 160M
@@
As shown in Table 1, we evaluate our approach on both short (16-frame) and long (64-frame) generation tasks. For 64-frame generation, our full model pushes FVD to 721.6. For 16-frame, we significantly reduce computational costs, achieving 6.7× faster inference (7.2s vs 48s) compared to CogVideo. This efficiency advantage becomes more pronounced in 64-frame generation, where our approach maintains 5.5× faster inference speed. Our staged training demonstrates clear benefits: Stage1 (trained with coarse annotations and $\mathcal{L}_{base}$ only) achieves a FVD of 482.1, which reduces to 438.3 with fine-grained annotations, and further reduces to 418.9 with $\mathcal{L}_{flow}$.
@@ A.2 Why is it Natural for GRID to Leverage Built-in Attention Mechanism
A.3 Comparison with Existing Approaches and Computational Efficiency Analysis
Current approaches to video generation can be categorized into two distinct paradigms, each with fundamental limitations in terms of architectural design and computational requirements. We provide a detailed analysis of these approaches and contrast them with our method:
Paradigm 1: Image Models as Single-Frame Generators Methods like SVD and AnimateDiff utilize pre-trained text-to-image models as frame generators while introducing separate modules for motion learning. This approach presents several fundamental limitations:
First, these methods require complex architectural additions for temporal modeling, introducing significant parameter overhead without leveraging the inherent capabilities of pre-trained image models. For instance, AnimateDiff introduces temporal attention layers that must be trained from scratch, while SVD requires separate motion estimation networks.
Second, the sequential nature of frame generation in these approaches leads to substantial computational overhead during inference. This sequential processing not only impacts generation speed but also limits the model’s ability to maintain long-term temporal consistency, as each frame is generated with limited context from previous frames.
Paradigm 2: End-to-End Video Architectures Recent approaches like Sora, CogVideo, and Huanyuan Video attempt to solve video generation through end-to-end training of video-specific architectures. While theoretically promising, these methods face severe practical constraints:
The computational requirements are particularly striking:
  • CogVideo requires approximately 35M video clips and an additional 2B filtered images from LAION-5B and COYO-700M datasets
  • Open-Sora necessitates more than 35M videos for training
  • These models typically demand multiple 80GB GPUs with sequence parallelism just for inference
  • Training typically requires thousands of GPU-days, making reproduction and iteration challenging for most research teams
Our Grid-based Framework: A Resource-Efficient Alternative In contrast, GRID achieves competitive performance through a fundamentally different approach:
1. Architectural Efficiency: Our grid-based framework requires only 160M additional parameters while maintaining competitive performance. This efficiency stems from:
  • Treating temporal sequences as spatial layouts, enabling parallel processing
  • Leveraging existing image generation capabilities without architectural complexity
  • Efficient parameter sharing across temporal and spatial dimensions
2. Data Efficiency: We achieve remarkable data efficiency improvements:
@@
$\text{Data Reduction}\approx\frac{>35M\text{ videos (previous methods)}}{<35K\text{ videos (our method)}}=1000\times$ (11)
This efficiency is achieved through:
  • Strategic use of grid-based training that maximizes information extraction from each video
  • Effective transfer learning from pre-trained image models
  • Focused training on essential video-specific components
@@
3. Computational Accessibility: Our approach enables high-quality video generation while maintaining accessibility for research environments with limited computational resources:
  • Training can be completed on standard research GPUs
  • Inference requires significantly less memory compared to end-to-end approaches
  • The model maintains strong performance across both video and image tasks
This comprehensive analysis demonstrates that our approach not only addresses the limitations of existing methods but also achieves substantial improvements in computational efficiency while maintaining competitive performance. The significant reductions in data requirements and computational resources make our method particularly valuable for practical applications and research environments with limited resources.
A.4 Distinction from In-Context LoRA
Recent work IC-LoRA (Huang et al., 2024b, a) also utilizes grid-based layouts for image generation, which might superficially appear similar to our approach. However, a careful analysis reveals fundamental differences in both theoretical foundation and technical implementation.
Different Theoretical Foundations: The core principle of IC-LoRA is to use grid layouts as a prompt engineering technique, where multiple images are arranged together to provide in-context examples for task adaptation. This is essentially an extension of in-context learning from language models to visual domain. Their grid layout serves merely as a presentation format for example-based learning.
In contrast, our approach fundamentally re-conceptualizes temporal sequences into spatial layouts. Rather than using grids for example presentation, we treat them as an inherent representation of temporal information, where spatial relationships in the grid directly correspond to temporal relationships in the sequence. This enables our model to learn and generate temporal dynamics in a holistic manner.
Distinct Technical Objectives: IC-LoRA’s technical implementation focuses on task adaptation through example pairs. Their method relies on LoRA-based fine-tuning and natural language prompts to define relationships between grid elements. However, this approach has inherent limitations in handling temporal dynamics, as it treats each grid element independently without explicit modeling of their temporal relationships.
Our method, on the other hand, is specifically designed for temporal sequence generation. We introduce parallel flow-matching and dedicated temporal loss functions that explicitly model motion patterns and temporal coherence. This allows our approach to capture and generate complex temporal dynamics that are beyond the capability of example-based methods like IC-LoRA.
Different Application Scopes: While IC-LoRA excels at static, example-based generation tasks through prompt engineering, it struggles with temporal sequence generation due to its fundamental design limitations. Our method, specifically designed for temporal modeling, naturally handles both static and dynamic visual generation tasks while maintaining precise control over temporal dynamics.
This analysis demonstrates that despite the superficial similarity in using grid layouts, our approach represents a fundamentally different direction in visual generation. We independently developed our method to address the specific challenges of temporal sequence generation, resulting in distinct technical contributions that go beyond the capabilities of example-based frameworks like IC-LoRA.
These crucial differences are evidenced by our method’s superior performance in temporal tasks and its ability to maintain consistent motion patterns across sequences - capabilities that are fundamentally beyond the scope of IC-LoRA’s example-based approach.
A.5 Inference Details
@@
A.6 Post-Processing Pipeline
@@ A.7 Potential Applications
A.7.1 Creative Multi-view Generation
@@
A.7.2 Flexible Frame Extension
@@
A.7.3 Future Extension to Video Understanding
@@
A.7.4 Maintained Image Generation Ability
@@ A.8 Limitations
A.9 Multyview Camera Parameters